13-7.VIP-MD is a health maintenance organization (HMO) located in North Carolina. Unlike the traditional fee-for-service model that determines the payment according to the actual services used or costs incurred, VIP-MD receives a fixed, prepaid amount from subscribers. The per member, per month rate (PMPM) is determined by estimating the health care cost per enrollee within a geographic location. The average health care coverage in North Carolina costs $371 per month, which is the same amount irrespective of the subscriber’s age. Because individuals are demanding quality care at reasonable rates, VIP-MD must contain its costs to remain competitive. A major competitor, National Physicians, entered the North Carolina market early in the current year with a monthly premium of $328. VIP-MD wants to maintain its current market penetration and hopes to increase its enrollees in the current year. The latest data on the number of enrollees and the associated costs follow:

Age Enrollment in Current Year Projected Enrollment Next Year Average Monthly Cost in Current Year
1–4 45,988 49,277 11,148,172
5–14 82,756 84,963 10,059,932
15–19 96,173 96,187 8,437,124
20–24 66,546 68,182 9,539,724
25–34 133,796 132,854 26,432,508
35–44 167,176 175,746 38,882,408
45–54 85,796 91,189 22,742,236
55–64 99,924 102,223 28,692,012
65–74 156,588 161,859 49,518,444
75–84 68,195 72,765 33,433,060
85 years and older 23,799 27,149 24,286,775
Total 1,026,737 1,062,394 263,172,395

Required:

1. Calculate the target cost required for VIP-MD to maintain its current market share and profit per enrollee in the current year.

2. Costs in the health care industry applicable to VIP-MD and National Physicians are expected to increase by 7% in the coming year. VIP-MD is planning for the year ahead and is expecting all providers, including VIP-MD and National Physicians, to increase their rates by $25 to $353. Calculate the new target cost assuming again that VIP-MD wants to maintain the same profit per enrollee as in the current year.
Answer #1

A.Cost per enrollee of current year = 263172395/1026737= $256.32

Since the cost per enrolle is to stay the same next year,

Estimated cost = cost per enrollee*expected enrollee=

256.32*1062394= $272312830.08

B. Cost to be increased by 7% next year= 256.32+7%= $263.32

New estimated cost = 263.32*1062394 = $279749588.08